# Linear realization of SU(3) parity doublet model for octet baryons with bad diquark

### v1

@@ Abstract
We construct a linear $SU(3)_{L}\times SU(3)_{R}$ parity doublet model for octet baryons. Our model employs the $(3,\bar{3})+(\bar{3},3)$ and $(3,6)+(6,3)$ chiral representations while excluding the $(8,1)+(1,8)$ representation. Through systematic analysis, we demonstrate that the $(3,6)+(6,3)$ representation containing symmetric “bad” diquarks, despite being energetically disfavored, is essential for reproducing the correct baryon mass hierarchy—particularly the $\Sigma$-$\Xi$ mass ordering. The model incorporates both spontaneous and explicit chiral symmetry breaking, with the latter implemented through bare quark mass terms that properly account for $SU(3)$ flavor breaking effects. Our numerical analysis successfully reproduces the ground-state octet baryon masses and predicts the spectrum of excited states up to 2.5 GeV. For the experimentally challenging $\Xi$ sector, we provide specific predictions for spin-parity assignments: identifying $\Xi(1950)$ as the first positive-parity excitation. The analysis reveals that ground states are dominated by the $(3,\bar{3})+(\bar{3},3)$ representation, consistent with the preference for “good” diquark configurations, while the $(3,6)+(6,3)$ contribution remains crucial for the mass spectrum.
@@ I Introduction
The study of baryon masses and their chiral structure remains one of the fundamental challenges in hadron physics. While the traditional view attributes baryon masses primarily to spontaneous chiral symmetry breaking through quark condensates, recent lattice QCD studies have revealed the existence of chiral-invariant mass contributions that persist even at the chiral restoration point [1, 2, 3]. This has motivated the development of parity doublet models (PDM) that incorporate both chiral-variant and chiral-invariant mass terms [4, 5, 6, 7, 8, 9].
@@
In contrast, the linear realization of chiral symmetry, where chiral partners are explicitly included as dynamical degrees of freedom, offers a more suitable framework for studying physics near and beyond the chiral restoration scale [21, 22, 23, 24, 25, 26, 8, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. However, despite its theoretical advantages, relatively little work has been done on $SU(3)_{L}\times SU(3)_{R}$ linear sigma models with parity doublet structure for baryons. The increased complexity arising from the numerous possible chiral representations and the need to properly implement both spontaneous and explicit symmetry breaking has made systematic studies challenging [31, 42, 43].
@@ IV $SU(3)$ Parity doublet structure
$\mathcal{L}_{0}=\mathcal{L}_{{\rm kin}}+\mathcal{L}_{{\rm CIM}}+\mathcal{L}_{{\rm Self}}+\mathcal{L}_{{\rm Inter}}$ (54)
@@ V.1 Brief review of Gell-Mann-Okubo mass relation
$\frac{{\rm tr}(M_{N})+{\rm tr}(M_{\Xi})}{2}=\frac{3,{\rm tr}(M_{\Lambda})+{\rm tr}(M_{\Sigma})}{4},$ (85)
@@ V.2 Numerical fitting
Our model shows good agreement with experimental data across all baryon sectors, and the GMO relation is satisfied for both the ground and excited states. In the nucleon sector, both the ground state $N(939)$ and the Roper resonance $N(1440)$ are well reproduced. The negative-parity $N(1535)$ appears slightly below its experimental value, which is consistent with expectations given that this state likely has meson-baryon molecular structure not captured in our approach. For the $\Lambda$ baryons, the model accurately reproduces the ground state $\Lambda(1116)$, first positive-parity excitation $\Lambda(1600)$ and the first negative-parity excitation $\Lambda(1670)$. Additionally, we predict a second negative-parity state to have the mass from 1800 MeV to 2500 MeV, which we identify with the $\Lambda(1800)$ resonance observed experimentally. For $\Sigma$ sector, our model successfully describes both the ground state $\Sigma(1193)$ and the first positive-parity excitation $\Sigma(1660)$. We identify two negative-parity states corresponding to the experimentally observed $\Sigma(1750)$ and $\Sigma(1900)$ resonances.
@@
where the coefficients satisfy the normalization condition $c_{\psi_{1}}^{2}+c_{\eta_{1}}^{2}+c_{\psi_{2}}^{2}+c_{\eta_{2}}^{2}=1$. Figure 6 displays the fraction of the $(3,\bar{3})+(\bar{3},3)$ representation (quantified by $c_{\psi_{1}}^{2}+c_{\psi_{2}}^{2}$) in the ground-state baryons as a function of the chiral invariant masses $(m_{0}^{(1)},m_{0}^{(2)})$. The uncolored regions correspond to parameter choices where $f_{\rm min}^{\rm best}>1$, indicating poor fits to the experimental data. First, we find that to have a good fit ($f_{\rm min}^{\rm best}\leq 1$) requires both chiral invariant masses to be sufficiently large. Specifically, $m_{0}^{(1)}$ must exceed 600 MeV, consistent with constraints from neutron star observations in the SU(2) parity doublet model [36, 57, 58, 59, 60, 61, 62]. Similarly, $m_{0}^{(2)}$ must lie within the range 650–1250 MeV, in agreement with our previous studies. Second, across all successful parameter regions, the $(3,\bar{3})+(\bar{3},3)$ representation consistently dominates the ground-state composition, with $c_{\psi_{1}}^{2}+c_{\psi_{2}}^{2}>0.5$. This dominance becomes even more pronounced as $m_{0}^{(2)}$ increases for fixed $m_{0}^{(1)}$, reflecting the physical expectation that the “bad” diquark configurations in the $(3,6)+(6,3)$ representation are energetically suppressed.
@@ VI Summary and Discussion
In this work, we have developed a linear $SU(3)_{L}\times SU(3)_{R}$ parity doublet model that successfully describes the octet baryon mass spectrum. Our approach makes several key contributions to understanding baryon structure within the chiral symmetry framework.
First, we demonstrated through systematic analysis why specific chiral representations are necessary for reproducing the correct baryon mass hierarchy. Starting from the minimal $(3,\bar{3})+(\bar{3},3)$ representation alone, we showed that this leads to unphysical degeneracies in the mass spectrum. While adding the $(8,1)+(1,8)$ representation resolves these degeneracies, it introduces incorrect mass orderings. This analysis establishes that the $(3,6)+(6,3)$ representation, despite containing energetically disfavored “bad” diquarks, is essential for achieving the correct mass spectrum—particularly the $\Sigma$-$\Xi$ mass ordering.
Second, our simplified model excludes the $(8,1)+(1,8)$ representation, reducing the number of parameters while retaining the essential physics. This choice is motivated by both theoretical considerations—the $(8,1)$ representation lacks clear diquark-quark separation and cannot be constructed from standard $SU(6)$ wave functions (See additional discussion in Appendix A).
Third, we incorporated explicit chiral symmetry breaking through quark mass terms, which was absent in previous analyses[42, 43]. This allows us to properly account for $SU(3)$ flavor breaking effects and study the interplay between spontaneous and explicit symmetry breaking. Our analysis reveals that flavor symmetry breaking from different meson condensates dominates over explicit chiral symmetry breaking from quark mass differences.
@@
Finally, we note that our calculation yields axial charge $g_{A}\approx 0.1$, significantly smaller than the experimental value $g_{A}^{\exp}=1.27$. This suppression arises from the substantial mixing between the original fields and their mirror partners, required by the large chiral-invariant masses ($m_{0}^{(1)}>600$ MeV and $m_{0}^{(2)}$ in the range 650–1250 MeV) necessary for successful fits. As demonstrated in Appendix B for the SU(2) parity doublet case, increasing chiral-invariant masses enhances mixing between chiral partners, which systematically reduces the axial charge—a feature that persists in our SU(3) framework. While this discrepancy appears problematic, it can be resolved by including higher-order derivative terms that contribute to the axial charge without affecting the mass spectrum. These derivative couplings, which we have not included in this leading-order analysis provide additional contribution to the axial charge. In Appendix C we also presents the extended Goldberger-Treiman relation connecting the axial charges to pion-nucleon couplings in our $SU(3)$ parity doublet model.

### v2

@@ Abstract
We construct a linear $SU(3)_{L}\times SU(3)_{R}$ parity doublet model for octet baryons. Our model employs the $(3,\bar{3})+(\bar{3},3)$ and $(3,6)+(6,3)$ chiral representations while excluding the $(8,1)+(1,8)$ representation. Through systematic analysis, we demonstrate that the $(3,6)+(6,3)$ representation containing symmetric “bad” diquarks, despite being energetically disfavored, is essential for reproducing the correct baryon mass hierarchy—particularly the $\Sigma$-$\Xi$ mass ordering. The model incorporates both spontaneous and explicit chiral symmetry breaking, with the latter implemented through bare quark mass terms that properly account for $SU(3)$ flavor breaking effects. Our numerical analysis successfully reproduces the ground-state octet baryon masses and predicts the spectrum of excited states up to 2.5 GeV. For the experimentally challenging $\Xi$ sector, we identify $\Xi(1950)$ as the first positive-parity excitation. The analysis reveals that ground states are dominated by the $(3,\bar{3})+(\bar{3},3)$ representation, consistent with the preference for “good” diquark configurations, while the $(3,6)+(6,3)$ contribution remains crucial for the mass spectrum.
@@ I Introduction
The study of baryon masses and their chiral structure remains one of the fundamental challenges in hadron physics. While the traditional view attributes baryon masses primarily to spontaneous chiral symmetry breaking through quark condensates, theoretical frameworks based on parity doublet structure that incorporate both chiral-variant and chiral-invariant mass terms have been developed to provide a more complete description of baryon properties [1, 2, 3, 4, 5, 6]. This theoretical framework has received strong support from recent lattice QCD studies, which have revealed the existence of chiral-invariant mass contributions that persist even at the chiral restoration point [7, 8, 9]. Furthermore, lattice QCD calculations of axial charges for negative-parity nucleon resonances have shown that $N^{*}(1535)$ has a remarkably small axial charge $g_{A}\sim\mathcal{O}(0.1)$, while $N^{*}(1650)$ exhibits $g_{A}\sim 0.5$, providing important insights into the chiral structure of excited baryons [10].
@@
In contrast, the linear realization of chiral symmetry, where chiral partners are explicitly included as dynamical degrees of freedom, offers a more suitable framework for studying physics near and beyond the chiral restoration scale [22, 23, 24, 25, 26, 27, 5, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. However, despite its theoretical advantages, relatively little work has been done on $SU(3)_{L}\times SU(3)_{R}$ linear sigma models with parity doublet structure for baryons. The increased complexity arising from the numerous possible chiral representations and the need to properly implement both spontaneous and explicit symmetry breaking has made systematic studies challenging [32, 43, 44, 45].
@@ IV $SU(3)$ Parity doublet structure
$\mathcal{L}_{0}=\mathcal{L}_{{\rm kin}}+\mathcal{L}_{{\rm CIM}}+\mathcal{L}_{{\rm diag}}+\mathcal{L}_{{\rm off-diag}}$ (54)
@@ V.1 Brief review of Gell-Mann-Okubo mass relation
$\frac{{\rm tr}(M_{N})+{\rm tr}(M_{\Xi})}{2}=\frac{3{\rm tr}(M_{\Lambda})+{\rm tr}(M_{\Sigma})}{4},$ (85)
@@ V.2 Numerical fitting
Our model shows good agreement with experimental data across all baryon sectors, and the GMO relation is satisfied for both the ground and excited states. In the nucleon sector, both the ground state $N(939)$ and the Roper resonance $N(1440)$ are well reproduced. The negative-parity $N(1535)$ appears slightly below its experimental value, which is consistent with expectations given that this state likely has meson-baryon molecular structure not captured in our approach [53, 54, 55]. For the $\Lambda$ baryons, the model accurately reproduces the ground state $\Lambda(1116)$, first positive-parity excitation $\Lambda(1600)$ and the first negative-parity excitation $\Lambda(1670)$. Additionally, we predict a second negative-parity state to have the mass from 1800 MeV to 2500 MeV, which we identify with the $\Lambda(1800)$ resonance observed experimentally. For $\Sigma$ sector, our model successfully describes both the ground state $\Sigma(1193)$ and the first positive-parity excitation $\Sigma(1660)$. We identify two negative-parity states corresponding to the experimentally observed $\Sigma(1750)$ and $\Sigma(1900)$ resonances.
@@
where the coefficients satisfy the normalization condition $c_{\psi_{1}}^{2}+c_{\eta_{1}}^{2}+c_{\psi_{2}}^{2}+c_{\eta_{2}}^{2}=1$. Figure 6 displays the fraction of the $(3,\bar{3})+(\bar{3},3)$ representation (quantified by $c_{\psi_{1}}^{2}+c_{\psi_{2}}^{2}$) in the ground-state baryons as a function of the chiral invariant masses $(m_{0}^{(1)},m_{0}^{(2)})$. The uncolored regions correspond to parameter choices where $f_{\rm min}^{\rm best}>1$, indicating poor fits to the experimental data. First, we find that to have a good fit ($f_{\rm min}^{\rm best}\leq 1$) requires both chiral invariant masses to be sufficiently large. Specifically, $m_{0}^{(1)}$ must exceed 600 MeV, consistent with constraints from neutron star observations in the SU(2) parity doublet model [37, 62, 63, 64, 65, 66, 67, 68]. Similarly, $m_{0}^{(2)}$ must lie within the range 650–1250 MeV, in agreement with our previous studies. Second, across all successful parameter regions, the $(3,\bar{3})+(\bar{3},3)$ representation consistently dominates the ground-state composition, with $c_{\psi_{1}}^{2}+c_{\psi_{2}}^{2}>0.5$. This dominance becomes even more pronounced as $m_{0}^{(2)}$ increases for fixed $m_{0}^{(1)}$, reflecting the physical expectation that the “bad” diquark configurations in the $(3,6)+(6,3)$ representation are energetically suppressed.
@@ VI Summary and Discussion
In this work, we have developed a linear $SU(3)_{L}\times SU(3)_{R}$ parity doublet model that successfully describes the octet baryon mass spectrum. In the context of SU(2) chiral symmetry, Ref. [69] proposed a quartet scheme where $\Delta_{\pm}$ and $N^{*}_{\pm}$ with even and odd parity form a chiral multiplet in the $(1,\frac{1}{2})\oplus(\frac{1}{2},1)$ representation. Using the mirror assignment, they derived parameter-free mass relations among the quartet members, which were shown to be well satisfied by observed resonances with spin $J=\frac{1}{2},\frac{3}{2},\frac{5}{2}$. Their work demonstrated that chiral symmetry provides meaningful constraints on excited baryon spectra and established the theoretical foundation for studying parity doubling beyond the ground-state nucleon. The present work extends this approach to the three-flavor sector.
Through systematic analysis, we demonstrated why specific chiral representations are necessary for reproducing the correct baryon mass hierarchy. Starting from the minimal $(3,\bar{3})+(\bar{3},3)$ representation alone, we showed that this leads to unphysical degeneracies in the mass spectrum. While adding the $(8,1)+(1,8)$ representation resolves these degeneracies, it introduces incorrect mass orderings. This analysis establishes that the $(3,6)+(6,3)$ representation, despite containing energetically disfavored “bad” diquarks, is essential for achieving the correct mass spectrum—particularly the $\Sigma$-$\Xi$ mass ordering. Our simplified model excludes the $(8,1)+(1,8)$ representation, reducing the number of parameters while retaining the essential physics. This choice is motivated by theoretical considerations: the $(8,1)$ representation lacks clear diquark-quark separation and cannot be constructed from standard $SU(6)$ wave functions (see additional discussion in Appendix A).
We also incorporated explicit chiral symmetry breaking through quark mass terms, which was absent in previous analyses [43, 44]. This allows us to properly account for $SU(3)$ flavor breaking effects and study the interplay between spontaneous and explicit symmetry breaking. Our analysis reveals that flavor symmetry breaking from different meson condensates dominates over explicit chiral symmetry breaking from quark mass differences.
@@
Finally, we note that our calculation yields axial charge $g_{A}\approx 0.1$, significantly smaller than the experimental value $g_{A}^{\exp}=1.27$. This suppression arises from the substantial mixing between the original fields and their mirror partners, required by the large chiral-invariant masses ($m_{0}^{(1)}>600$ MeV and $m_{0}^{(2)}$ in the range 650–1250 MeV) necessary for successful fits. As demonstrated in Appendix B for the SU(2) parity doublet case, increasing chiral-invariant masses enhances mixing between chiral partners, which systematically reduces the axial charge—a feature that persists in our SU(3) framework. While this discrepancy appears problematic, it can be resolved by including higher-order derivative terms that contribute to the axial charge without affecting the mass spectrum. These derivative couplings, which we have not included in this leading-order analysis provide additional contribution to the axial charge [70, 71] while leaving the static baryon masses unchanged. In Appendix C we also presents the extended Goldberger-Treiman relation connecting the axial charges to pion-nucleon couplings in our $SU(3)$ parity doublet model.
Several important extensions of this work is possible for future investigation. First, the suppressed axial charge $g_{A}\approx 0.1$ can be remedied by incorporating higher-order derivative couplings, which contribute to the axial current without modifying the mass spectrum [70, 71]. With realistic $g_{A}$ values in hand, a systematic calculation of meson-baryon coupling constants and decay widths becomes feasible through the extended Goldberger-Treiman relation derived in Appendix C. This would enable quantitative predictions for resonance widths, particularly for the Roper $N(1440)$ and the hyperon excited states, providing further tests of our model against experimental data. Second, a proper treatment of the $\eta$ meson and $\eta^{\prime}$ meson requires incorporating $U_{A}(1)$ anomaly effects, which would allow us to study the $^{*}(1535)-NN$-$\eta$ coupling and address the strong affinity of $N^{*}(1535)$ for the $\eta N$ channel. Third, the connection between our mirror assignment and the five-quark (pentaquark) structure suggests that explicit pentaquark degrees of freedom might play a role in describing states with significant meson-baryon molecular components. Finally, extending the present framework to finite density and temperature would enable applications to neutron star matter and heavy-ion collisions, where chiral symmetry restoration through parity doubling is expected to play a crucial role.